Convert the rectangular point (-8,8sqrt3) to polar coordinates with r>0.
Also give two additional ways to represent this point in polar coordinates
for which r<0 and -2pi *and* r>0 and -2pi
Answer by stanbon(48558)

Convert the rectangular point (-8,8sqrt3) to polar coordinates with r>0.
r=sqrt(8^2+(8sqrt3)^2)= sqrt(256)= 16
theta = tan^-1((8sqrt3)/-8)=tan^-1(-sqrt3)=120 degrees or (2/3)pi
polar form: (16,(2/3)pi)
-----------------------
(-16,(5/3)pi)

Also give two additional ways to represent this point in polar coordinates
for which r<0 and -2pi (-16,(-pi/3))
------------------
*and* r>0 and -2pi (16,((-4/3)pi))
=============
Cheers,
Stan H.